A block of mass 5 kg is kept on a rough horizontal
surface (μ = 0.5). The block can be pulled
horizontally by an unknown force. The contact
force between the surface and block can be : (Force
is acting horizontally)
(A) 50 N (B) 25 N
(C) 75 N (D) 52 N

Grade:12

3 Answers

vikas askiitian expert
13 years ago

normal  reaction = mg = 5*10 = 50N

option A) should be correct..

ramendra bahadur singh
13 years ago

as the mass is to be pulled upward so the minimum force required is equal to the weight of body                    w=mg        =>           w=F     F=5*9.8=49N        OR          F=5*10=50N        HENCE OPTION (A) IS CORRECT

sourabh singh
13 years ago

75